# The unknotting number, hard unknot diagrams, and reinforcement learning

### v1

@@ 3.3. Braids
When considering diagrams of connected sums, one can obtain potentially better mixing of the components by inserting up to 7 braid words equivalent to the identity, and which mixes strands between the two components, compared to overlaying. See, for example, Figure 6.
@@
We furthermore trained an RL agent which operated directly on braid words rather than using invariants. In this case, we used a transformer architecture [37], a ML architecture designed to work on sequences. The input to the model is a sequence of integers representing the braid word, and the output is – akin to the invariant-based agent – a probability distribution over which crossing to switch. This had the big advantage of being invariant-free and hence very fast. It performed well on smaller braids ($\leq 60$ crossings, 3–8 strands), but struggled to unknot larger braids efficiently. Future work could investigate invariant-free unknotting agents further.
@@ 4. Additivity of the unknotting number
Figure 7. The distribution of random knots $K$ in our dataset by SnapPy simplified crossing number and by $u(K)$.
To construct $S$, we obtained $u(K)$ for 31k random knots between 10 and 60 crossings and 26k quasipositive knots betwen 10 and 50 crossings in their SnapPy simplified diagram with large $|\sigma(K)|$, and where the upper bounds given by the RL agent and the lower bounds from knot Floer homology coincided. We found it difficult to generate random knots $K$ with $|\sigma(K)|$ large. See Figure 7 for the distribution of the crossing number of a SnapPy simplified diagram and $u(K)$ for the random knots in our dataset. For the lower bounds, we used the signature, the Ozsváth–Szabó $\tau$ invariant, and $|\nu(\pm K)|$, which is sometimes 1 bigger than $|\tau(K)|$. We also added torus knots and at most 12-crossing knots with known unknotting number from KnotInfo.
@@
In order to construct a non-trivial diagram of $K\#K^{\prime}$, we overlaid the diagrams of $K$ and $K^{\prime}$; see Figure 8. Note that, in an overlay sum, every crossing arc only intersects the connected sum sphere at most once. Hence, in approximately 100k cases, we also performed random Reidemeister moves to further mix the components. We did not always do this as the limit of our RL agent was around 200 crossings. For some experiments, we also considered connected overlay sums of 3 at most 12-crossing knots, with crossing number up to 120. In addition, we also searched among connected sums obtained from braids by inserting identity braid words, as explained in Section 3.3. See Figure 9 for the performance of the RL agent based on the Jones polynomial on connected sums, where the $x$-axis shows the sum of the unknotting numbers of the summands.
Figure 9. The performance of the RL agent based on the Jones polynomial on connected sums and no forced inter-component crossing changes.
@@ 4.1. Strong conjecture
We have found 20 counterexamples to Conjecture 1.2 by starting from diagrams where the RL agent found unknotting sets with several hyperbolic inter-component crossings and performing all in-component crossing changes from the unknotting set. These produce connected sums where the inter-component crossing changes result in a diagram of the unknot. One such example is shown in Figure 8, where the inter-component crossings 13, 14, 48, and 50 are unknotting.
@@
The following result was proved by Lickorish for tangles. His proof extends immediately to generalised tangles, and is therefore omitted.
@@
For $i\in\{1,2\}$, isotope a sub-arc $a_{i}$ of $K_{i}$ such that it becomes straight and parallel to the $y$-axis in $\mathbb{R}^{3}$. Furthermore, let $C_{i}$ be a set of disjoint unknotting framed crossing arcs of $K_{i}$ such that $|C_{i}|=u(K_{i})$, and such that they are disjoint from $a_{i}$. Isotope the initial point $c(0)$ of each arc $c\in C_{i}$ into $a_{i}$. Then contract $c$ to a straight line segment via an ambient isotopy of $\mathbb{R}^{3}$ fixing $a_{i}$ such that $z(c(0))>z(c(1))$, where $z\colon\mathbb{R}^{3}\to\mathbb{R}$ is the $z$-coordinate function (this can be achieved by performing a finger move on $K_{i}$ by moving $c(1)\in K_{i}$ along $c$). We finally shorten and move the vertical arcs in $C_{i}$ so close to each other such that the corresponding crossings in the diagram $\mathcal{D}_{i}$ obtained by perturbing $K_{i}$ and projecting it onto the $(x,y)$-plane are consecutive along the projection of $a_{i}$. Let $p_{i}\in a_{i}$ be a point such that $y(p_{i})<y(c(0))$ for every $c\in C_{i}$. We take the connected sum of $K_{1}$ and $K_{2}$ at $p_{1}$ and $p_{2}$. We denote the resulting diagram of $K_{1}\#K_{2}$ by $\mathcal{D}$. See the upper left of Figure 10. In $K_{1}\#K_{2}$, the arcs $a_{1}$ and $a_{2}$ will become arcs $a_{+}$ and $a_{-}$, where $a_{+}$ is the upper and $a_{-}$ the lower horizontal strand in the upper left of Figure 10. Note that $a_{+}$ contains the crossings $C_{1}\cup C_{2}$.
@@
First, suppose that $u(K_{1})>1$ and $u(K_{2})>1$. Let $c_{i}\in C_{i}$ be the crossing closest to $p_{i}$. Then perform a finger move on the lower strand at $c_{1}$ along $a_{+}$, isotoping it across the crossings in $C_{2}$. Next, isotope the lower strand at $c_{2}$ along $a_{+}$, isotoping it across the crossings in $C_{1}$. The resulting diagram $\mathcal{D}^{\prime}$ is shown in the upper right of Figure 10. We write $c_{i}^{\prime}$ for the crossing of $\mathcal{D}^{\prime}$ corresponding to $c_{i}$. Changing all the crossings of $\mathcal{D}$ and $\mathcal{D}^{\prime}$ along $a_{+}$ lead to equivalent diagrams, so, to the unknot. Let us write
(1) $C_{i}^{\prime}:=C_{i}\setminus\{c_{i}\}\cup\{c_{i}^{\prime}\}.$
If we change a single crossing in $C_{2}^{\prime}$, after an isotopy, the diagram looks like the bottom left of Figure 10 near the connected sum sphere. Analogously, changing a crossing in $C_{1}^{\prime}$ leads to the diagram in the bottom right of Figure 10.
We claim that if we change any crossing $c\in C_{1}^{\prime}\cup C_{2}^{\prime}$, the resulting knot $K$ is prime. Without loss of generality, suppose that $c\in C_{2}^{\prime}$, so we have the situation shown in the lower left of Figure 10. The knot $K$ is split as a sum of two tangles $T_{1}$ and $T_{2}$ by the connected sum sphere $S$ of $K_{1}\#K_{2}$, where $T_{i}$ lies on the same side of $S$ as $K_{i}$, for $i\in\{1,2\}$.
@@
We claim that $T_{1}$ contains no non-trivial ball-arc pair. Suppose that it did. We have already observed that we can attach a trivial tangle to the outside of $T_{1}$ to obtain $K_{1}$. Hence, if $T_{1}$ did contain a non-trivial ball-arc pair, this would be knotted like $K_{1}$ as $K_{1}$ is prime. Let $K^{\prime}_{1}$ be the result of changing the crossing $c_{1}$ of $K_{1}$. As $c_{1}$ is part of a minimal unknotting set for $K_{1}$, we have $u(K^{\prime}_{1})=u(K_{1})-1$, so $K_{1}^{\prime}\neq K_{1}$. We note that we can attach a trivial tangle to the outside of $T_{1}$ and obtain the knot $K^{\prime}_{1}$. The non-trivial ball-arc pair in $T_{1}$ then forms a summand for $K^{\prime}_{1}$. Hence, $K^{\prime}_{1}=K_{1}\#J$ for some knot $J$, where $J\neq U$ since $K^{\prime}_{1}\neq K_{1}$. This contradicts the assumption that $K^{\prime}_{1}$ is not the connected sum of $K_{1}$ and a non-trivial knot. Thus, we have shown that $T_{1}$ is a non-trivial tangle that contains no non-trivial ball-arc pair, and is hence prime.
@@
We claim that $A$ is prime. It is certainly not a trivial tangle, since it has a simple closed curve component. Suppose that there is a 2-sphere $S$ intersecting the tangle $A$ transversely in two points that does not bound an unknotted arc. Form $A\cup A$, glued with a rotation through $\pi/2$. This has a prime non-split alternating diagram, and hence, by a theorem of Menasco [25], it is a prime non-split link. But $S$ would force this link either to be split or not prime, which is a contradiction, proving the claim.
@@
We are now in a position to show that $T_{2}$ has no non-trivial ball-arc pair. Suppose that, on the contrary, there is a sphere $S^{\prime}$ in $T_{2}$ such that $|S^{\prime}\pitchfork T_{2}|=2$ and which bounds a non-trivial arc. This arc is therefore knotted like a non-trivial knot $K^{\prime}$. Then $|S^{\prime}\pitchfork L^{\prime}_{2}|=2$, since $L^{\prime}_{2}$ is the union of $T_{2}$ and a trivial tangle. On one side, $S^{\prime}$ bounds an arc knotted like $K^{\prime}$. On the other side, it also does not bound a trivial arc, since $L^{\prime}_{2}$ is a link of two components. Hence, $L^{\prime}_{2}$ is not prime, contrary to what we proved above.
@@
Now suppose that $u(K_{1})=1$ and $u(K_{2})>1$. The case $u(K_{2})=1$ and $u(K_{1})>1$ is analogous. Let $c_{1}$ be the only crossing in $C_{1}$, and let $c_{2}$ be the crossing in $C_{2}$ furthest from $p_{2}$; see the left of Figure 12. As before, perform a finger move on the lower strand at $c_{1}$ across the crossings in $C_{2}$. We denote by $c_{1}^{\prime}$ the crossing corresponding to $c_{1}$; see the middle of Figure 12. Now perform a finger move on the lower strand at $c_{2}$ along $a_{+}$, moving it across $c_{1}^{\prime}$, and resulting in the new crossing $c_{2}^{\prime}$; see the right of Figure 12. Call this diagram $\mathcal{D}^{\prime}$, and let $C^{\prime}_{i}$ be as in equation (1) for $i\in\{1,2\}$. A proof similar to the above shows that, if we change any crossing $c\in C_{1}^{\prime}\cup C_{2}^{\prime}$, we obtain a prime knot $K$. The only difference happens when $c=c_{1}^{\prime}$, which we now discuss.
As before, we split $K$ into tangles $T_{1}$ and $T_{2}$. The same argument shows that $T_{1}$ is prime. To see that $T_{2}$ is non-trivial, note that we can glue a trivial tangle to it to obtain a link $L_{2}$, which is isotopic to $K_{2}$ with an unknot linking the crossing $c_{2}$ non-trivially; see Figure 13. If $T_{2}$ were trivial, then $L_{2}$ would be a 2-bridge link. Hence, $K_{2}$ would be a trivial knot, contradicting the assumption that it is prime.
We can again write $L_{2}$ as a sum of tangles $A$ and $B$, where $A$ is shown in Figure 11. We have already proved that $A$ is prime. Let $E$ be the trivial tangle obtained from $A$ by deleting the unknot component, and $E^{\prime}$ the trivial tangle obtained by changing the crossing $c_{2}^{\prime}$ of $E$. With this notation, the same argument as before shows that $B$ is also prime, and so $L_{2}=A\cup B$ is prime. It follows that $T_{2}$ is prime and hence that $K=T_{1}\cup T_{2}$ is prime. ∎
@@
$c_{1}$
$c_{1}^{\prime}$
$c_{2}$
$c_{2}^{\prime}$
Figure 12. In the proof of Theorem 1.3, when the summand $K_{1}$ has $u(K_{1})=1$, we perform the isotopies shown to obtain a diagram where every crossing change results in a prime knot.
$c_{2}^{\prime}$
Figure 13. The left shows the link $L_{2}$ obtained from the tangle $T_{2}$ by adding a trivial tangle. The right shows $L_{2}$ after the isotopy of the unknot component into a neighbourhood of the crossing $c_{2}$ of $K_{2}$.
@@ 4.2. New unknotting numbers assuming additivity of $u$
If we assume that $u$ is additive and consider knots that appear along length $u(K)+u(K^{\prime})$ unknotting trajectories of connected sums $K\#K^{\prime}$ where $u(K)$ and $u(K^{\prime})$ are both known, we obtain the unknotting number of 43 knots $K$ with $c(K)\leq 12$ that are unknown; see Table 1. More specifically, we considered knots that appear along minimal unknotting trajectories of counterexamples $\mathcal{D}$ to Conjecture 1.2 by changing $u(\mathcal{D})-4$ or $u(\mathcal{D})-5$ crossings in a pre-computed unknotting set of $\mathcal{D}$, then simplifying, obtaining a diagram $\mathcal{D}^{\prime}$, and then brute-forcing all unknotting trajectories from $\mathcal{D}^{\prime}$.
@@
Table 1. At most 12-crossings knots with unknown unknotting numbers that we found on minimal unknotting trajectories of connected sums $K\#K^{\prime}$ where $u(K)$ and $u(K^{\prime})$ are known.
In all these examples, $u(K)$ was equal to the KnotInfo upper bound. If one of these knots had lower unknotting number than the upper bound, it would imply that the unknotting number is not additive.
@@
Except for 12a898, 12a916, 12a981, and 12a999, all of these knots $K$ have a crossing change in their KnotInfo diagram $\mathcal{D}$ that results in a connected sum $K_{0}\#K_{1}$ with $u(K_{0})$ and $u(K_{1})$ known and

### v2

@@ 3.3. Braids
Let $\beta_{1}$ and $\beta_{2}$ be braids of indices $k_{1}$ and $k_{2}$, respectively. Furthermore, let $\sigma_{k_{1}}$ be the generator of the braid group of index $k_{1}+k_{2}$ that swaps strands $k_{1}$ and $k_{1}+1$. Then $\widehat{\beta_{1}\sigma_{k_{1}}\beta_{2}}=\widehat{\beta}_{1}\#\widehat{\beta}_{2}$. We further mixed $\widehat{\beta_{1}\sigma_{k_{1}}\beta_{2}}$ by inserting braid words that are equivalent to the identity, and which mix strands between the two summands. We chose 7 such braid words to keep the size of the resulting braids manageable by the RL agent. See, for example, Figure 5. This method seemed to result in different mixing between the connected summands than overlaying them (see Section 4 for a discussion on overlaying diagrams).
@@
We furthermore trained an RL agent which operated directly on braid words rather than using invariants. In this case, we used a transformer architecture [37], an ML architecture designed to work on sequences. The input to the model is a sequence of integers representing the braid word, and the output is – akin to the invariant-based agent – a probability distribution over which crossing to switch. This had the big advantage of being invariant-free and hence very fast. It performed well on smaller braids ($\leq 60$ crossings, 3–8 strands), but struggled to unknot larger braids efficiently. Future work could investigate invariant-free unknotting agents further.
@@ 4. Additivity of the unknotting number
Figure 6. The distribution of random knots $K$ in our dataset by SnapPy simplified crossing number, by $u(K)$, and by $\sigma(K)$.
To construct $S$, we obtained $u(K)$ for 31k random knots between 10 and 60 crossings and 26k quasipositive knots betwen 10 and 50 crossings in their SnapPy simplified diagram with large $|\sigma(K)|$, and where the upper bounds given by the RL agent and the lower bounds from knot Floer homology coincided. We found it difficult to generate random knots $K$ with $|\sigma(K)|$ large. See Figure 6 for the distribution of the crossing number of a SnapPy simplified diagram and the unknotting numbers and signatures for the random knots in our dataset. For the lower bounds, we used the signature, the Ozsváth–Szabó $\tau$ invariant, and $|\nu(\pm K)|$, which is sometimes 1 bigger than $|\tau(K)|$. (Note that $\sigma$ and $\tau$ are symmetric with respect to mirroring, while $\nu$ is not.) We also added torus knots and at most 12-crossing knots with known unknotting number from KnotInfo.
@@
In order to construct a non-trivial diagram of $K\#-K^{\prime}$, we overlaid the diagrams of $K$ and $K^{\prime}$; see Figure 7. Note that, in an overlay sum, every crossing arc only intersects the connected sum sphere at most once. Hence, in approximately 100k cases, we also performed random Reidemeister moves to further mix the components. We did not always do this as the limit of our RL agent was around 200 crossings. For some experiments, we also considered connected overlay sums of 3 at most 12-crossing knots, with crossing number up to 120. In addition, we also searched among connected sums obtained from braids by inserting identity braid words, as explained in Section 3.3. See Figure 8 for the performance of the RL agent based on the Jones polynomial on connected sums, where the $x$-axis shows the sum of the unknotting numbers of the summands.
Figure 8. The performance of the RL agent based on the Jones polynomial on connected sums and no forced inter-component crossing changes. The four dashed lines indicate a difference of 3, 6, 9, and 12, respectively, between the diagram unknotting number found by the agent and the unknotting number expected assuming additivity.
@@ 4.1. Strong conjecture
We have found 20 counterexamples to Conjecture 1.2 by starting from diagrams where the RL agent found unknotting sets with several hyperbolic inter-component crossings and performing all in-component crossing changes from the unknotting set. These produce connected sums where the remaining crossing changes result in a diagram of the unknot. In general, some of the inter-component crossings may become in-component after this procedure, but we have checked this is not the case for these 20 examples. One such example is shown in Figure 7, where the inter-component crossings 13, 14, 48, and 50 are unknotting.
@@
The following result was proved by Lickorish [22] for tangles. His proof extends immediately to generalised tangles, and is therefore omitted.
@@
For $i\in\{1,2\}$, isotope a sub-arc $a_{i}$ of $K_{i}$ such that it becomes straight and parallel to the $y$-axis in $\mathbb{R}^{3}$; see Figure 9. Furthermore, let $C_{i}$ be a set of disjoint unknotting framed crossing arcs of $K_{i}$ such that $|C_{i}|=u(K_{i})$, and such that they are disjoint from $a_{i}$. Isotope the initial point $c(0)$ of each arc $c\in C_{i}$ into $a_{i}$. Then contract $c$ to a straight line segment via an ambient isotopy of $\mathbb{R}^{3}$ fixing $a_{i}$ such that $z(c(0))>z(c(1))$, where $z\colon\mathbb{R}^{3}\to\mathbb{R}$ is the $z$-coordinate function (this can be achieved by performing a finger move on $K_{i}$ by moving $c(1)\in K_{i}$ along $c$). We finally shorten and move the vertical arcs in $C_{i}$ so close to each other such that the corresponding crossings in the diagram $\mathcal{D}_{i}$ obtained by perturbing $K_{i}$ and projecting it onto the $(x,y)$-plane are consecutive along the projection of $a_{i}$. Let $p_{i}\in a_{i}$ be a point such that $y(p_{i})<y(c(0))$ for every $c\in C_{i}$. We take the connected sum of $K_{1}$ and $K_{2}$ at $p_{1}$ and $p_{2}$. We denote the resulting diagram of $K_{1}\#K_{2}$ by $\mathcal{D}$. See the upper left of Figure 10. In $K_{1}\#K_{2}$, the arcs $a_{1}$ and $a_{2}$ will become arcs $a_{+}$ and $a_{-}$, where $a_{+}$ is the upper and $a_{-}$ the lower horizontal strand in the upper left of Figure 10. From now on, by a slight abuse of notation, $C_{1}$ and $C_{2}$ will denote sets of crossings rather than crossing arcs. Note that $a_{+}$ contains the crossings $C_{1}\cup C_{2}$.
$a_{2}$
$a_{1}$
$K_{1}$
$C_{1}$
$C_{2}$
$K_{2}$
$p_{1}$
$p_{2}$
Figure 9. The knots $K_{i}$ with the vertical arcs $a_{i}$, unknotting set of crossings $C_{i}$, and connected sum points $p_{i}$ for $i\in\{1,2\}$.
@@
$c_{1}^{\prime}$
$d_{2}^{\prime}$
Figure 11. The result of changing the crossing $c_{1}^{\prime}$.
Suppose that $u(K_{2})>1$. The case $u(K_{1})>1$ is analogous. For $i\in\{1,2\}$, let $c_{i}\in C_{i}$ be the crossing closest to $p_{i}$ and let $d_{2}$ be the crossing in $C_{2}$ furthest from $p_{2}$. Then perform a finger move on the lower strand at $c_{1}$ along $a_{+}$, isotoping it across the crossings in $C_{2}$. Next, isotope the lower strand at $c_{2}$ along $a_{+}$, isotoping it across the crossings in $C_{1}$. We write $c_{i}^{\prime}$ for the crossing corresponding to $c_{i}$. Perform a finger move on the lower strand at $d_{2}$ along $a_{+}$, moving it across $c_{1}^{\prime}$, and resulting in the new crossing $d_{2}^{\prime}$. The resulting diagram $\mathcal{D}^{\prime}$ is shown in the upper right of Figure 10. Changing all the crossings of $\mathcal{D}$ and $\mathcal{D}^{\prime}$ along $a_{+}$ lead to equivalent diagrams, so, to the unknot. Let us write
(1) $C_{1}^{\prime}:=(C_{1}\setminus\{c_{1}\})\cup\{c_{1}^{\prime}\}\text{ and }C_{2}^{\prime}:=(C_{2}\setminus\{c_{2},d_{2}\})\cup\{c_{2}^{\prime},d_{2}^{\prime}\},$
and set $C:=C_{1}^{\prime}\cup C_{2}^{\prime}$.
If we change a single crossing in $C_{2}^{\prime}$, after an isotopy, the diagram looks like the bottom left of Figure 10 near the connected sum sphere. Analogously, changing a crossing in $C_{1}^{\prime}\setminus\{c_{1}^{\prime}\}$ leads to the diagram in the bottom right of Figure 10. Finally, changing $c_{1}^{\prime}$, after an isotopy, results in the situation shown in Figure 11.
We claim that if we change any crossing $c\in C$, the resulting knot $K$ is prime. First, assume that $c\neq c_{1}^{\prime}$. Without loss of generality, suppose that $c\in C_{2}^{\prime}$, so, after an isotopy, we have the situation shown in the lower left of Figure 10 (this is also the case for $c=c_{2}^{\prime}$). The knot $K$ is split as a sum of two tangles $T_{1}$ and $T_{2}$ by the connected sum sphere $S$ of $K_{1}\#K_{2}$, where $T_{i}$ lies on the same side of $S$ as $K_{i}$, for $i\in\{1,2\}$.
@@
We claim that $T_{1}$ contains no non-trivial ball-arc pair. Suppose that it did. We have already observed that we can attach a trivial tangle to the outside of $T_{1}$ to obtain $K_{1}$. Hence, if $T_{1}$ did contain a non-trivial ball-arc pair, this would be knotted like $K_{1}$ as $K_{1}$ is prime. Let $K^{\prime}_{1}$ be the result of changing the crossing $c_{1}$ of $K_{1}$. As $c_{1}$ is part of a minimal unknotting set for $K_{1}$, we have $u(K^{\prime}_{1})=u(K_{1})-1$, so $K_{1}^{\prime}\neq K_{1}$. We note that we can also attach a trivial tangle in a different way to the outside of $T_{1}$ and obtain the knot $K^{\prime}_{1}$. The non-trivial ball-arc pair in $T_{1}$ then forms a summand for $K^{\prime}_{1}$. Hence, $K^{\prime}_{1}=K_{1}\#J$ for some knot $J$, where $J\neq U$ since $K^{\prime}_{1}\neq K_{1}$. This contradicts the assumption that $K^{\prime}_{1}$ is not the connected sum of $K_{1}$ and a non-trivial knot. Thus, we have shown that $T_{1}$ is a non-trivial tangle that contains no non-trivial ball-arc pair, and is hence prime.
@@
We claim that $A$ is prime. It is certainly not a trivial tangle, since it has a simple closed curve component. Suppose that there is a 2-sphere $S^{\prime}$ intersecting the tangle $A$ transversely in two points that does not bound an unknotted arc. Glue two copies of $A$ with a rotation through $\pi/2$. This has a prime non-split alternating diagram, and hence, by a theorem of Menasco [25], it is a prime non-split link. But $S^{\prime}$ would force this link either to be split or not prime, which is a contradiction, proving the claim.
@@
We are now in a position to show that $T_{2}$ has no non-trivial ball-arc pair. Suppose that, on the contrary, there is a sphere $S^{\prime\prime}$ in $T_{2}$ that intersects $T_{2}$ in two points and which bounds a non-trivial arc. This arc is therefore knotted like a non-trivial knot $K^{\prime}$. Then $S^{\prime\prime}$ transversely intersects $L_{2}^{\prime}$ in two points, since $L^{\prime}_{2}$ is the union of $T_{2}$ and a trivial tangle. On one side, $S^{\prime\prime}$ bounds an arc knotted like $K^{\prime}$. On the other side, it also does not bound a trivial arc, since $L^{\prime}_{2}$ is a link of two components. Hence, $L^{\prime}_{2}$ is not prime, contrary to what we proved above.
@@
We now consider the case when $c=c_{1}^{\prime}$; see Figure 11. As before, we split $K$ into tangles $T_{1}$ and $T_{2}$. The same argument shows that $T_{1}$ is prime. To see that $T_{2}$ is non-trivial, note that we can glue a trivial tangle to it to obtain a link $L_{2}$, which is isotopic to $K_{2}$ with an unknot linking the crossing $d_{2}$ non-trivially; see Figure 13. If $T_{2}$ were trivial, then $L_{2}$ would be a 2-bridge link. Hence, $K_{2}$ would be a trivial knot, contradicting the assumption that it is prime.
We can again write $L_{2}$ as a sum of tangles $A$ and $B$, where $A$ is shown in Figure 12. We have already proved that $A$ is prime. Let $E$ be the trivial tangle obtained from $A$ by deleting the unknot component, and $E^{\prime}$ the trivial tangle obtained by changing the crossing $d_{2}^{\prime}$ of $E$. With this notation, the same argument as before shows that $B$ is also prime, and so $L_{2}=A\cup B$ is prime. It follows that $T_{2}$ is prime and hence that $K=T_{1}\cup T_{2}$ is prime. ∎
@@
$d_{2}^{\prime}$
Figure 13. The left shows the link $L_{2}$ obtained from the tangle $T_{2}$ by adding a trivial tangle. The right shows $L_{2}$ after the isotopy of the unknot component into a neighbourhood of the crossing $d_{2}$ of $K_{2}$.
@@ 4.2. New unknotting numbers assuming additivity of $u$
If we assume that $u$ is additive and consider knots that appear along length $u(K)+u(K^{\prime})$ unknotting trajectories of connected sums $K\#K^{\prime}$ where $u(K)$ and $u(K^{\prime})$ are both known, we obtain the unknotting number of 43 knots $K^{\prime\prime}$ with $c(K^{\prime\prime})\leq 12$ that are unknown; see Table 1. More specifically, we considered knots that appear along minimal unknotting trajectories of counterexamples $\mathcal{D}$ to Conjecture 1.2 by changing $u(\mathcal{D})-4$ or $u(\mathcal{D})-5$ crossings in a pre-computed unknotting set of $\mathcal{D}$, then simplifying, obtaining a diagram $\mathcal{D}^{\prime}$, and then brute-forcing all unknotting trajectories from $\mathcal{D}^{\prime}$.
@@
Table 1. At most 12-crossings knots with unknown unknotting numbers that we found on minimal (assuming the additivity of the unknotting number) unknotting trajectories of connected sums $K\#K^{\prime}$ where $u(K)$ and $u(K^{\prime})$ are known.
In all these examples, $u(K^{\prime\prime})$ was equal to the KnotInfo upper bound. If one of these knots had lower unknotting number than the upper bound, it would imply that the unknotting number is not additive.
@@
Except for 12a898, 12a916, 12a981, and 12a999, all of these knots $K^{\prime\prime}$ have a crossing change in their KnotInfo diagram $\mathcal{D}$ that results in a connected sum $K_{0}\#K_{1}$ with $u(K_{0})$ and $u(K_{1})$ known and